# DyCodeEval: Dynamic Benchmarking of Reasoning Capabilities in Code Large Language Models Under Data Contamination

### v1

@@ Abstract
The rapid evolution of code largelanguage models (LLMs) underscores the need for effective and transparent benchmarking of their reasoning capabilities. However, the current benchmarking approach heavily depends on publicly available, human-created datasets. The widespread use of these fixed benchmark datasets makes the benchmarking process to be static and thus particularly susceptible to data contamination—an unavoidable consequence of the extensive data collection processes used to train Code LLMs. Existing approaches that address data contamination often suffer from human effort limitations and imbalanced problem complexity. To tackle these challenges, we propose DyCodeEval, a novel benchmarking suite for evaluating Code LLMs under potential data contamination. Given a seed programming problem, DyCodeEval employs multiple agents to extract and modify the context without altering the core logic, generating semantically equivalent variations. We introduce a dynamic data generation methods and conduct empirical studies on two seed datasets across 21 Code LLMs. Results show that DyCodeEval effectively benchmarks reasoning capabilities under contamination risks while generating diverse problem sets to ensure consistent and reliable evaluations. Our project webpage can be found at this link111https://codekaleidoscope.github.io/dycodeeval.html.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant potential as assistant software developers, particularly in code generation (Chen et al., 2021; Guo et al., 2024; Jiang et al., 2024; Di et al., 2024). Consequently, numerous code-focused LLMs have been developed. These models are trained on vast corpora of natural language and programming language data. Once well-trained, they can comprehend human instructions and generate corresponding code snippets.
@@
As diverse model architectures and training algorithms for code LLMs continue to emerge (Vaswani et al., 2017; Shazeer et al., 2017), a key focus in code LLM research is the effective benchmarking of each model’s code reasoning capability. For instance, each model is realeased on HuggingFace every week. Without a standardized and transparent benchmarking suite, assessing these models’ performance and driving improvements becomes a significant challenge.
@@
However, existing benchmarking suites for evaluating code LLMs are inadequate due to their static benchmarking schema, which can lead to potential data contamination from unintended data crawling. Research suggests that such contamination may already be present in current LLMs (Brown et al., 2020; Jain et al., 2024). Although some methods aim to provide contamination-free benchmarking for code LLMs, they still rely on manual efforts. For example, LiveCodeBench (Jain et al., 2024) proposes crawling new programming problems from online platforms and benchmarking LLMs based on timestamps, while PPM (Chen et al., 2024) attempts to systematize new programming problems by combining manually defined operators. However, these methods have several limitations: (1) Significant Manual Effort: These methods still require substantial manual input to create such datasets. For example, PPM necessitates manually defining the lambda operator, while LiveCodeBench shifts the burden of manual design to question authors on coding platforms. (2) Imbalanced Semantic Complexity: The newly generated benchmarking datasets often lack semantic equivalence with the original ones. As a result, when a model performs worse on these benchmarks, it is challenging to determine whether the lower score reflects diminished model capabilities or increased benchmark complexity. And thus, these new benchmark’s results fail to provide meaningful guidance for model developers to improve their models effectively.
@@
To address this limitation, rather than manually creating benchmarking datasets with uncertain semantic complexity, we aim to develop an automated method for dynamically evaluating code LLMs. However, designing such a method presents two key challenges: (1) Generating Semantically Diverse Yet Complexity-Controlled Problems. The first challenge is that how to ensure the generated problems vary in semantics while maintaining controlled complexity; and (2) Providing Comprehensive Benchmarking. A proper benchmark programming problem must include fine-grained test cases and canonical solutions to rigorously assess correctness.
@@
The performance of Code LLMs on our dynamic benchmarks degraded significantly, suggesting potential data contamination (§4.3).
@@ 2 Background & Related Work
2.1 Benchmarking LLMs
Numerous benchmarks have appeared in order to test the effectiveness of LLMs (Hendrycks et al., 2021; Clark et al., 2018; Cobbe et al., 2021). Many emphasize an understanding of knowledge and language such as Glue (Wang et al., 2019) and SciQ (Welbl et al., 2017). Others focus on reasoning capabilities e.g. DROP (Dua et al., 2019) and BBH (Suzgun et al., 2022). And some focus on specific capabilities such as coding (Chen et al., 2021). Fig. 1 shows an example of a programming problem from a popular coding benchmark HumanEval, where a programming problem usually includes three parts: prompt, canonical solution, and test cases.
2.2 Data Contamination Free Benchmarking LLMs
Data contamination has become a prevalent issue in benchmarking LLMs (Brown et al., 2020; Jain et al., 2024). Many researchers have developed methods in hopes of mitigating its effect. Given the public nature of benchmarks, one set of approaches (Jacovi et al., 2023; Rajore et al., 2024) aims to implement ideas such as encryption and privatization in order to protect the data from training use. Other attempts to solve this problem include DyVal (Zhu et al., 2024a) which leverages the structure of DAGs to dynamically generate evaluation sets, TreeEval (Li et al., 2024) which uses a high performing LLM as an examiner utilizing a tree planning strategy, and ITD (Inference Time Decontamination) (Zhu et al., 2024c) which detects and rewrites leaked samples of benchmarks without changing its complexity. Unfortunately, the limitation of these existing methods is the lack of complexity equivalence with the original problem set. Therefore, there is a need to devise a benchmark that maintains equivalence while providing contamination free data.
2.3 LLM as Judgement Agent
@@
Recently, LLMs have become increasingly used as examiners given their capabilities of analyzing large amounts of data and providing unbiased assessments (Bai et al., 2023; Fernandes et al., 2023). This growing trend has gained interest for two reasons: (1) Enhanced generation of training/testing data (Li et al., 2024; Liu et al., 2024) (2) Accurate evaluation and comparison of LLM outputs such as in PandaLM (Wang et al., 2024) and DyVal (Zhu et al., 2024b). Additionally, as LLMs have been able to perform remarkbly well on unseen tasks, they offer a faster, equally accurate alternative to human evaluation, (Chiang & yi Lee, 2023).
@@ 3.1 Design Overview
The design overview of DyCodeEval is shown in Fig. 2. Given a seed programming problem from existing benchmarks, DyCodeEval generates a semantically different yet complexity-equivalent problem using a metamorphic relationship. DyCodeEval comprises four agents: (1) Scenario Proposer, (2) Context Generator, (3) Prompt Rewriter, and (4) Validator. The Scenario Proposer suggests real-world domains (e.g., banking, healthcare, education), from which DyCodeEval randomly selects one. The Context Generator then analyzes input types in the canonical solution and assigns a relevant context for each input variable based on the selected scenario. The Prompt Rewriter reformulates the problem to align with the input variable contexts and chosen scenario. Finally, the Validator ensures the new problem remains consistent with the original. If inconsistencies are detected, DyCodeEval will repeat the aforementioned process until a valid variant is produced.
@@ 3.2 Detailed Design
Validation Agent. Although we provide the LLM with detailed scenario and context information for rewriting, there are cases where the rewriting agent unintentionally alters the consistency. To address this, we design a validation agent to assess whether the generated question maintains the integrity of the original intent and informational content. The validation prompt is designed from two angles: (1) it directs the LLM to compare the seed programming problem prompt with the rephrased prompt, ensuring the preservation of the core concept and factual accuracy, and (2) it asks the LLM to check whether the seed canonical solutions align with the generated programming problem prompt. Specifically, we design two comparison prompts to query the LLM and retain only those rewritten prompts for which both comparison responses are ”YES”.
@@ 3.3 Theoretical Collision Analysis
DyCodeEval generates programming problems dynamically with randomness, reducing the risk of potential data contamination. To analyze this, we conduct a collision analysis. The randomness in DyCodeEval arises from both the scenario proposal and context generation phases. We assume the scenario proposer generates scenarios, and for each scenario, the context generation produces contexts, while ignoring randomness in the rewriting phase. Based on this, we present the following theorem.
@@
After running DyCodeEval times on the same seed problem, Then the probability that the samples after the first one are all different from the first sampled item satisfies: .
@@
Consider the seed dataset of size , After running DyCodeEval times on this dataset, If , Then the probability that the generated dataset after the first one are all not the same as the first generated dataset satisfies:
The proof could be found in Appendix A
@@ 4.2 Benchmarking Contaminated Model
Models. We conduct our study with three public-available Code LLMs: Llama-3.2-1B, Llama-3.2-3B, and DeepSeek-Coder-1.3b. The selected code generation models are diverse in terms of model architecture, model size, and training methods.
Model Contamination Process. For each model, we simulate data contamination by intentionally leaking a portion of the benchmarking dataset during fine-tuning. (1) For the static benchmarking method, we directly include part of the benchmarking dataset in fine-tuning. (2) For our method, we run it twice on each seed dataset to generate two versions of new programming problems, one for training and the other for benchmarking. We experiment with leaked data percentages of 0%, 25%, 50%, 75%, and 100%, producing four distinct contaminated models. Each polluted model is then evaluated on the benchmarking dataset using the Pass@1 metric. The formal definition of Pass@1 is shown in (1), where is the number of the generated solution candidate, and is the number of the correct solutions that can pass all test cases.
@@
From the results, we make the following observations: (1). Data contamination creates a false sense of code reasoning capability under static benchmarks. When the benchmarking dataset is leaked and used for fine-tuning, the model achieves a higher Pass@1 score on the corresponding benchmark. However, this improvement does not accurately reflect the model’s true reasoning ability, as its performance declines on other benchmarks that were not included in fine-tuning. (2). Our dynamic benchmarking mitigates the impact of data contamination. Different from static benchmarks, our approach prevents contaminated models from achieving artificially high Pass@1 scores after fine-tuning. This is due to the randomness in our method, which ensures minimal or no overlap between different runs, reducing the risk of direct data leakage. (3). Our dynamic benchmarking dataset provides results comparable to manually curated, non-contaminated datasets. In static benchmarking, as the percentage of leaked data increases, the model’s Pass@1 score on the contaminated benchmark steadily improves. However, its performance on other benchmarks remains relatively stable, showing little variation across different contamination levels. Interestingly, this stability also applies to our method. If the base model is not contaminated on the selected seed dataset, this suggests that our approach provides competitive benchmarking results similar to those of human-curated datasets. (4). A notable anomaly is observed in DeepSeek-Coder. When only 25% of the benchmarking dataset is used for fine-tuning, the model’s Pass@1 score drops below that of the original, unmodified model. We hypothesize that the model may already be overfitted to the contaminated dataset, and further fine-tuning with limited data could destabilize this overfitting without providing enough new information to help the model adapt.
4.3 Benchmarking On-the-wild Model
@@
We then apply DyCodeEval to benchmark more on-the-wild code LLMs, besides the model used in §4.2. We consider the following code LLMs: Llama-3.1-8B, CodeLlama-7b, CodeLlama-13b, DeepSeek-V2-Lite, DeepSeek-Coder-V2-Lite-Base, Llama-3.1-8B-Instruct, Qwen2.5-Coder-7B, Qwen2.5-7B-Instruct, Qwen2.5-7B, Claude-3.5-haiku, Claude-3.5-sonnet,Qwen2.5-Coder-7B-Instruct .
@@
The results are presented in Fig. 5, with the left figure showing the results on HumanEval and the right showing the results on MBPP. In each figure, the x-axis represents the Pass@1 scores on our generated dataset, and the y-axis represents the Pass@1 scores on the seed dataset. The blue region corresponds to the regression area of the on-the-wild model, the red region represents the regression area of the overfitted model on this dataset, and the orange area indicates the overfitted model on the other dataset.
@@
From these results, we observe that for both seed datasets, the on-the-wild model’s Pass@1 scores maintain a linear relationship, while the overfitted model appears as an outlier. A notable finding from our on-the-wild evaluation is that the model Qwen2.5-Coder-7B consistently falls outside the 95% confidence interval of the regression area, suggesting it may be contaminated on both datasets.
4.4 Programming Problem Diversity

### v2

@@ Abstract
The rapid advancement of code large language models (Code LLMs) underscores the critical need for effective and transparent benchmarking methods. However, current benchmarking predominantly relies on publicly available, human-created datasets. The widespread use of these static benchmark datasets makes the evaluation process particularly susceptible to data contamination—an unavoidable consequence of the extensive data collection processes employed during LLM training. Existing methods for addressing data contamination typically face significant limitations, including reliance on substantial human effort and difficulty in managing class imbalances. To overcome these challenges, we propose DyCodeEval, a novel benchmarking suite specifically designed to evaluate Code LLMs under realistic contamination scenarios. Given an initial seed programming problem, DyCodeEval utilizes multiple agents to systematically extract and modify contextual information without changing the core logic, generating semantically equivalent variations. We introduce a dynamic data generation method and conduct extensive empirical studies on two seed datasets involving 18 Code LLMs. The results demonstrate that DyCodeEval effectively assesses the reasoning capabilities of Code LLMs under contamination conditions while producing diverse problem variants, thereby ensuring robust and consistent benchmarking outcomes. Our project webpage can be found at this link111https://codekaleidoscope.github.io/dycodeeval.html.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant potential as assistant software developers, particularly in code generation (Chen et al., 2021; Guo et al., 2024; Jiang et al., 2024; Di et al., 2024). Consequently, numerous code-focused LLMs have been developed. These models are trained on vast corpora of natural language and programming language data. Once well trained, they can comprehend human instructions and generate the corresponding code snippets.
@@
As diverse model architectures and training algorithms for code LLMs continue to emerge (Vaswani et al., 2017; Shazeer et al., 2017), a key focus in code LLM research is the effective benchmarking of each model’s code reasoning capability. Without a standardized and transparent benchmarking suite, assessing these models’ performance and driving improvements becomes a significant challenge.
@@
However, existing benchmarking suites for evaluating code LLMs are inadequate due to their static benchmarking schema, which can lead to potential data contamination from unintended data crawling. Research suggests that such contamination may already be present in current LLMs (Chen et al., 2025; Jain et al., 2024; Dong et al., 2024). Although some methods aim to provide contamination-free benchmarking for code LLMs, they still rely on manual efforts. For example, LiveCodeBench (Jain et al., 2024) proposes crawling new programming problems from online platforms and benchmarking LLMs based on timestamps, while PPM (Chen et al., 2024) attempts to systematize new programming problems by combining manually defined operators. However, these methods have several limitations: (1) Significant Manual Effort: These methods still require substantial manual input to create such datasets. For example, PPM necessitates manually defining the lambda operator, while LiveCodeBench shifts the burden of manual design to question authors on coding platforms. (2) Imbalanced Semantic Complexity: The newly generated benchmarking datasets often lack semantic equivalence with the original ones. As a result, when a model performs worse on these benchmarks, it is challenging to determine whether the lower score reflects diminished model capabilities or increased benchmark complexity. Thus, these new benchmark results fail to provide meaningful guidance for model developers to improve their models effectively.
@@
To address this limitation, rather than manually creating benchmarking datasets with uncertain semantic complexity, we aim to develop an automated method for dynamically evaluating code LLMs. However, designing such a method presents two key challenges: (1) Generating Semantically Diverse Yet Complexity-Controlled Problems. The first challenge is how to ensure the generated problems vary in semantics while maintaining controlled complexity. (2) Providing Comprehensive Benchmarking. A proper benchmark programming problem must include fine-grained test cases and canonical solutions to rigorously assess correctness.
@@
The performance of some Code LLMs on our dynamic benchmarks degraded significantly, suggesting potential data contamination of these Code LLMs (§4.3).
@@ 2 Background & Related Work
2.1 Benchmarking Code LLMs
Code LLMs have been widely adopted in various real-world software engineering applications, leading to the development of numerous benchmarks for evaluating their capabilities in code understanding and reasoning (Chen et al., 2021; Li et al., 2024a; Guan et al., 2025; Austin et al., 2021; Chen et al., 2024; Austin et al., 2021; Yu et al., 2024; Jimenez et al., 2024; Ding et al., 2023; Mathai et al., 2024). Among the many tasks designed to assess code reasoning, this work focuses specifically on the task of natural language to code generation and reviews representative benchmarks in this area. HumanEval (Chen et al., 2021) introduced a human-crafted dataset to evaluate the code generation capabilities of large language models. EvalPlus (Liu et al., 2023) later identified the limitations of HumanEval and MBPP—particularly their limited test case coverage—and proposed a more rigorous benchmark. HumanEval-XL (Peng et al., 2024) further extended HumanEval to support multilingual settings. Fig. 1 illustrates an example from HumanEval, a widely used benchmark for the natural language to code generation task. Each programming problem typically consists of three components: a prompt, a canonical solution, and a set of test cases. The prompt is first fed into the Code LLM to generate a candidate solution, which is then executed against hidden test cases to evaluate its correctness.
2.2 Data Contamination Free Benchmarking
Data contamination has become a significant concern in benchmarking large language models (LLMs) (Brown et al., 2020; Jain et al., 2024; Chen et al., 2025), as it can lead to inflated performance scores and unreliable evaluations. To mitigate this issue, researchers have proposed various contamination-free benchmarking strategies, which can be broadly categorized into three approaches. The first line of work focuses on data protection through encryption and privatization. For instance, Jacovi et al.(Jacovi et al., 2023) and Rajore et al.(Rajore et al., 2024) propose techniques to safeguard benchmark data from being included in LLM training corpora. The second line of research emphasizes timely benchmark updates. LiveBench (White et al., 2024), for example, compiles questions from recent sources such as math competitions held within the past year and regularly updates its dataset. Similarly, LiveCodeBench (Jain et al., 2024) continuously collects new human-authored programming problems from online platforms like LeetCode to maintain freshness and reduce the risk of contamination. The third line of research explores dynamic generation of evaluation sets. DyVal (Zhu et al., 2024a) uses DAG structures to create dynamic benchmarks, TreeEval (Li et al., 2024b) employs high-performing LLMs to generate and evaluate problems via tree planning, and ITD (Inference-Time Decontamination) (Zhu et al., 2024c) identifies and rewrites leaked benchmark samples while preserving their complexity.
2.3 LLM as Judgment Agent
@@
Recently, LLMs have become increasingly used as examiners given their capabilities of analyzing large amounts of data and providing unbiased assessments (Bai et al., 2023; Fernandes et al., 2023). This growing trend has gained interest for two reasons: (1) Enhanced generation of training/testing data (Li et al., 2024b; Liu et al., 2024) (2) Accurate evaluation and comparison of LLM outputs such as in PandaLM (Wang et al., 2024) and DyVal (Zhu et al., 2024b). Additionally, as LLMs have been able to perform remarkbly well on unseen tasks, they offer a faster, equally accurate alternative to human evaluation, (Chiang & yi Lee, 2023).
@@ 3.1 Design Overview
The design overview of DyCodeEval is shown in Fig. 2. Given a seed programming problem from existing benchmarks, DyCodeEval generates a semantically different yet complexity-equivalent problem using a metamorphic relationship. DyCodeEval comprises of four agents: (1) Scenario Proposer, (2) Context Generator, (3) Prompt Rewriter, and (4) Validator. The Scenario Proposer suggests real-world domains (e.g., banking, healthcare, education) from which DyCodeEval randomly selects one. The Context Generator then analyzes input types in the canonical solution and assigns a relevant context for each input variable based on the selected scenario. The Prompt Rewriter reformulates the problem to align with the input variable contexts and chosen scenario. Finally, the Validator ensures the new problem remains consistent with the original. If inconsistencies are detected, DyCodeEval will repeat the aforementioned process until a valid variant is produced.
@@ 3.2 Detailed Design
Validation Agent. Although we provide the LLM with detailed scenario and context information for rewriting, there are cases where the rewriting agent unintentionally alters the consistency. To address this, we design a validation agent to assess whether the generated question maintains the integrity of the original intent and informational content. The validation prompt is designed from two angles: (1) it directs the LLM to compare the seed programming problem prompt with the rephrased prompt, ensuring the preservation of the core concept and factual accuracy, and (2) it asks the LLM to check whether the seed canonical solutions align with the generated programming problem prompt. Specifically, we design two comparison prompts to query the LLM and retain only those rewritten prompts for which both comparison responses are “YES”.
@@ 3.3 Theoretical Collision Analysis
DyCodeEval generates programming problems dynamically with randomness, reducing the risk of potential data contamination. To analyze this, we conduct a collision analysis. The randomness in DyCodeEval arises from both the scenario proposal and context generation phases. We assume the scenario proposer generates scenarios, and for each scenario, the context generation produces contexts, while ignoring randomness in the rewriting phase. We also assume that the random sampling process follows a uniform distribution. Based on this, we present the following theorem.
@@
After running DyCodeEval times on the same seed problem, then the probability that the samples after the first are all different from the first sampled item satisfies: .
@@
Consider the seed dataset of size , After running DyCodeEval times on this dataset, If , then the probability that the generated datasets after the first one are not the same as the first generated dataset satisfies:
The proof could be found in Appendix A.
@@ 4.2 Benchmarking Contaminated Model
Models. We conduct our study with three public-available Code LLMs: Llama-3.2-1B, Llama-3.2-3B, and DeepSeek-Coder-1.3b.
Model Contamination Process. For each model, we simulate data contamination by intentionally leaking a portion of the benchmarking dataset during fine-tuning. We experiment with leaked data percentages of 0%, 25%, 50%, 75%, and 100%, producing four distinct contaminated models. Each polluted model is then evaluated on the benchmarking dataset using the Pass@1 metric. The formal definition of Pass@1 is shown in (1), where is the number of the generated solution candidate, and is the number of the correct solutions that can pass all test cases.
@@
From the results, we make the following observations: (1) Data contamination creates a false sense of code reasoning capability under static benchmarks. When the benchmarking dataset is leaked and used for fine-tuning, the model achieves a higher Pass@1 score on the corresponding benchmark. However, this improvement does not accurately reflect the model’s true reasoning ability, as its performance declines on other benchmarks that were not included in fine-tuning. (2) Our dynamic benchmarking mitigates the impact of data contamination. Different from static benchmarks, our approach prevents contaminated models from achieving artificially high Pass@1 scores after fine-tuning. This is due to the randomness in our method, which ensures minimal or no overlap between different runs, reducing the risk of direct data leakage. (3) Our dynamic benchmarking dataset provides results comparable to manually curated, non-contaminated datasets. In static benchmarking, as the percentage of leaked data increases, the model’s Pass@1 score on the contaminated benchmark steadily improves. However, its performance on other benchmarks remains relatively stable, showing little variation across different contamination levels. Interestingly, this stability also applies to our method. If the base model is not contaminated on the selected seed dataset, this suggests that our approach provides competitive benchmarking results similar to those of human-curated datasets. (4) A notable anomaly is observed in DeepSeek-Coder. When only 25% of the benchmarking dataset is used for fine-tuning, the model’s Pass@1 score drops below that of the original, unmodified model. We hypothesize that the model may already be overfitted to the contaminated dataset, and further fine-tuning with limited data could destabilize this overfitting without providing enough new information to help the model adapt.
4.3 Benchmarking In-the-Wild Model
@@
We then apply DyCodeEval to benchmark more in-the-wild code LLMs, besides the models used in §4.2. We consider the following code LLMs: Llama-3.1-8B, CodeLlama-7b, CodeLlama-13b, DeepSeek-V2-Lite, DeepSeek-Coder-V2-Lite-Base, Llama-3.1-8B-Instruct, Qwen2.5-Coder-7B, Qwen2.5-7B-Instruct, Qwen2.5-7B, Claude-3.5-haiku, Claude-3.5-sonnet,Qwen2.5-Coder-7B-Instruct .
@@
The results are presented in Fig. 5, with the left figure showing the results on HumanEval and the right showing the results on MBPP. In each figure, the x-axis represents the Pass@1 scores on our generated dataset, and the y-axis represents the Pass@1 scores on the seed dataset. The blue region corresponds to the regression area of the in-the-wild model, the red region represents the regression area of the overfitted model on this dataset, and the orange area indicates the overfitted model on the other dataset.
@@
From these results, we observe that for both seed datasets, the in-the-wild model’s Pass@1 scores maintain a linear relationship, while the overfitted model appears as an outlier. A notable finding from our in-the-wild evaluation is that the model Qwen2.5-Coder-7B consistently falls outside the 95% confidence interval of the regression area, suggesting it may be contaminated on both datasets.
4.4 Problem Diversity
To evaluate the diversity of the generated programming problems, we conduct two experiments: one for external diversity and one for internal diversity. External diversity quantifies the dissimilarity between the generated and seed problems, while internal diversity measures the diversity within each problem-generation method across trials. We use two metrics: BLEU-4 to measure syntactical diversity and cosine similarity of the prompt’s semantic embedding to measure semantic diversity. For semantic embedding, we use the GPT-2 model to obtain the embedding of each natural language prompt. Moreover, we also consider PPM (Chen et al., 2024) and a series of robustness-based mutations (Wang et al., 2023), such as token replacement, insert blank lines, as our comparison baseline.